OpenStudy (anonymous):

HELP! If the simple interest on $3000 for 6 years is $1620, then what is the rate?

OpenStudy (anonymous):

I = PRT but we are looking for the rate, so let me change this formula to : I/PT = R I = 1620 P = 3000 T = 6 Now lets sub our stuff in the formula.. 1620/(3000)(6) = R 1620/18,000 = R .09 = R now we will multiply .09 by 100 to make it a percent and we get 9% 9% = R check... I = PRT 1620 = (3000)(.09)(6) 1620 = 1620 (correct)

OpenStudy (anonymous):

Perfect! Thank you!

OpenStudy (anonymous):

did you understand what I did or do you have any questions ?

OpenStudy (anonymous):

How did you know what formula to use?

OpenStudy (anonymous):

Are there other formulas I need to know?

OpenStudy (anonymous):

Simple interest formula is I = PRT. This formula can be rearranged if you are looking for something besides I. It can be changed to : I/PT = R I/RT = P

OpenStudy (anonymous):

Oh. OK! Thank you very much! That's helpful!

OpenStudy (anonymous):

Glad I could help.....now this is not the formula for compound interest, just simple interest

OpenStudy (anonymous):

Alrighty. Thanks!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

Ok. Question for you. How do I find r My teacher want me to use the formula for future value, A=P(1+rt) and elementary algebra to find the missing quantity. A=$1,920 ;P=$1,500 ;t= 4 years

OpenStudy (anonymous):

A = P(1 + rt) A = 1920 P = 1500 t = 4 now lets sub... 1920 = 1500(1 + 4r) -- distribute through the parenthesis 1920 = 1500 + 6000r -- subtract 1500 from both sides 1920 - 1500 = 6000r 420 = 6000r -- divide by 6000 420/6000 = r .07 = r If your teacher wants your r in a percent, multiply it by 100 and you get 7% check... 1920 = 1500(1 + (.07)(4) 1920 = 1500 (1 + .28) 1920 = 1500 (1.28) 1920 = 1920 (correct)